# Towards AI-Assisted Synthesis of Verified Dafny Methods

### v1

@@ 3. Study Design
  • RQ2 [Signature Prompting]: How does additional context (signatures and tests) affect the synthesis outcome of LLMs?
  • RQ3 [Dynamic Few-Shot Prompting]: Do prompts designed with Chain of Thought (CoT) and semantically similar few-shot examples explaining problem decomposition step by step improve the synthesis outcome of LLMs?
@@ 3.1. Test Dataset: MBPP-san-DFY
In order to conduct our study, we needed a collection of problems that has natural language problem description corresponding formally verified Dafny code. We searched Dafny code in GitHub that yielded only 188 repositories having no natural language specifications of the available code. We would have to recover the informal natural language specifications from the formal code but by doing so we might inject errors by using words with a different intention than the original developers, or inject our own biases.
Since we aim to synthesize Dafny code from natural language specifications , for this purpose, we adopted the Mostly Basic Programming Problems (MBPP) benchmark published by Google Research in 2021 (Austin et al., 2021), which was used in prior program synthesis and verification studies. MBPP consists of a collection of close to 1,000 crowd-sourced algorithmic tasks with text descriptions, function signatures, concise implementations in Python, and three test cases to check their functional correctness. The tasks focus on basic arithmetic, array and integer sequences, and string processing. The textual descriptions of these tasks are concrete enough that a human should be able to translate them into code without any additional clarifications.
Solutions to 427 MBPP tasks have been inspected manually, and published as a subset called MBPP-sanitized (MBPP-san). We based our data on MBPP-san. However, since our goal is to synthesize Dafny, we chose to remove tasks that relied on Python-specific data types or external libraries, such as regex, math, heapq, etc.  [^4] Dafny does not have a standard library, much less external libraries. Note that any Dafny libraries would need to be formally-verified a prior, or else the Dafny tasks would become much more complex than their Python versions. We culled 82 tasks, leaving a dataset of 345 problems.
In order to keep our study’s costs under budget, we randomly selected 228 problems. Out of these 228 problems, we picked a random sample of 50 to act as verified Dafny exemplars (see below Section 3.2) leaving us with 178 remaining problems for testing. We manually translated each of those 178 problem descriptions, signatures, and test cases into Dafny, but we did not translate the solutions’ Python code, or attempt to write specifications. This resulted in our main test dataset, MBPP-san-DFY: 178 instances of MBPP-san problems, each one with a description, method signature, and three test cases translated to suit Dafny.
The MBPP problem set, and Python solutions to the MBPP problems, are publicly available and it seems likely they would have been incorporated into the LLMs’ training sets. However, there are no Dafny implementations of MBPP. Therefore, we hypothesize that employing MBPP-san-DFY dataset would not create any data leakage in our experiment.
@@ 3.2. Human Written Dataset: MBPP-DFY-50
To perform Dynamic Few-Shot prompting, we needed a relatively large and varied collection of verified Dafny methods that could serve as examplars to be included in prompts. A search for an existing suitable collection proved fruitless. Our best alternative was to translate a subset of MBPP-san into Dafny ourselves, but now also translating the Python solution for each problems into Dafny, then coming up with specifications (method pre- and post-conditions, loop invariants) for each example, and finally doing whatever annotation, correction, or rewriting was needed until the Dafny verifier would signal that the method was correct.
@@
3.3. LLM Selection
Program synthesis in Dafny requires knowledge of Dafny syntax, as well as reasoning and theorem-proving abilities (Noble et al., 2022; First et al., 2023). Therefore, we looked for LLMs that possess a combination of these abilities. To determine which LLMs to use, we performed a pilot study using our handwritten dataset MBPP-DFY-50. We started by looking at GPT-4 (OpenAI, 2023a) and GPT-3.5 (Mark Chen et al., 2021) from Open-AI (OpenAI, 2023b), PaLM-2 (Rohan Anil et al., 2023) and PaLM (Aakanksha Chowdhery et al., 2022) from Google, and CodeT5+ (Wang et al., 2021), CodeT5 (Wang et al., 2023) from Salesforce and Code-Llama (Rozière et al., 2023) from Meta. All these LLMs have been explored in prior code generation, program synthesis, and reasoning-related studies. Our pilot study consisted of querying all these LLMs with the problem description, asking them to solve each. We then analyzed their responses to get a sense of their reasoning abilities, and the quality of their approaches compared to our ground truth. For these 50 problems, CodeT5, CodeT5+, LLaMA, Code-LLaMA, and PaLM were not able to synthesize a single syntactically correct Dafny method with specifications, as all of them encountered compilation errors. Even in GPT-3.5, we got only 5/50 syntactically correct solutions. However, GPT-4 and PaLM-2 generate verified Dafny methods in 16/50 and 7/50 problems respectively. Based on this, we decided to narrow our investigation to GPT-4 and PaLM-2 only.
@@ 3.4. Prompt Design
Based on the background literature and our pilot study, we knew that our experiment would depend on the prompts we offered to the LLMs. Therefore, we experimented with three different prompts.
3.4.1. RQ1 [Contextless]
To answer RQ1 we designed a simple prompt that identifies the programming language (Dafny), the problem description, and instructions for how to format the code response. Our prompt design is inspired by Austin et al. (Austin et al., 2021). This kind of prompt works well for popular languages like Python, JavaScript, Java, C#, and C++.
You are an expert Dafny programmer, and here is your task: Write a method in Dafny to find maximum run of uppercase characters in the given string. You must return the method in the following form: ```dafny //Dafny Code ```
You are an expert Dafny programmer, and here is your task: Write a method in Dafny to find maximum run of uppercase characters in the given string. Method Signature: method maxRunUppercase (str: string ) returns (res: int ) Your code should satisfy these tests: var out1:=maxRunUppercase( "GeMKSForGERksISBESt" ); assert out1==5; var out2:=maxRunUppercase( "PrECIOusMOVemENTSYT" ); assert out2==6; var out3:=maxRunUppercase( "GooGLEFluTTER" ); assert out3==4; You must return the method in the following form: ```dafny //Dafny Code ```
Figure 2. An example of Contextless and Signature prompt
@@ 3.4.1. RQ1 [Contextless]
The Contextless prompt instructs the LLMs to synthesize a Dafny method in a single step without providing any additional context. Figure 2 (top) shows an example of Contextless prompt. Since writing functional code in Dafny is similar to writing code in high-level programming languages, with Contextless prompt our goal is to establish a baseline of the LLMs’ knowledge of Dafny.
3.4.2. RQ2 [Signature]
@@
For RQ2, we extended the Contextless prompt with the desired method’s signature and three test cases (see Figure 2 (bottom)). Code generation experiments for non-verified code often include test cases and signatures (implicitly or explicitly) (Austin et al., 2021; Mark Chen et al., 2021). We hypothesized that with the method signature and test cases LLMs can infer the pre- and post-conditions that would be beneficial to synthesize complex problems.
3.4.3. RQ3 [Dynamic Few-Shot]
Work in LLMs has shown that Chain of Thought (CoT) reasoning to decompose a problem into multi-intermediate steps can lead to drastically improved reasoning, especially for code synthesis tasks (Madaan et al., 2022; Wei et al., 2022; Reynolds and McDonell, 2021; Beurer-Kellner et al., 2023).
Used together with relevant examples in the prompt, aka few-shot prompting, it becomes possible to “teach” LLMs things that had a weak presence in the training data. Inspired by this, we designed a prompt following the principle of Chain of Though (CoT) with semantically similar few-shot Dafny method examples where each example explains its problem decomposition step by step and how to derive the implementation in Dafny. The thought process and reasoning steps of this prompt was informed by our own experience with Dafny. When we write Dafny methods, we start by defining the method’s signature, followed by thinking about the pre- and post-conditions that should hold for that signature. Then we write the functional code, i.e. the algorithm that we believe should result in those postconditions to hold, and finally we add the proof, i.e. the necessary loop invariants and verification annotations that lead the Dafny verifier to conclude that the algorithm is, indeed, correct.
However, for a given problem, this prompt will be effective when semantically similar examples of the given problem are included in this prompt. Therefore, we sought for a heuristic that can be employed to search best examples for a given problem. While creating our MBPP-DFY-50 dataset, we observed that the postconditions are the most valuable parts of Dafny method specifications. For each problem we tackled, we searched through the problems we had already solved, looking for similar postconditions. We then used those examples to guide the proof of the new problem. We identified a pattern that problems having similar method signature and post-conditions require the same type of Dafny proofs and code to solve the problem. This pattern motivated us to develop a prompt chaining pipeline for selecting few-shot Dafny code examples that have similar method signatures and post-conditions for a given problem. Figure 3 depicts the high-level overview of our developed pipeline for prompt chaining.
@@ 3.4.3. RQ3 [Dynamic Few-Shot]
Figure 3. A high-level overview of the prompt-chaining pipeline for Dynamic Few-Shot Prompting.
@@
To select semantically similar Dafny methods based on postconditions, we employed prompt chaining in two phases. In the first phase, for a given problem, we simply prompt the LLM to generate the method specifications such as signature and postconditions. To teach LLM how to generate such specifications, we provided three few-shots examples of problem description, method signature and postcondition pairs from our MBPP-DFY-50 dataset. We elected these examples based on the semantic similarity of the given problem description. An example of such specification prompting is shown in Figure 4 (top).
@@
TASK: Write a method in Dafny to convert a given string to uppercase. Method Signature: method ToUppercase(s: string ) returns (v: string ) Postconditions: - The length of the returning string must be the same as the length of the input string - All lowercase characters of the input string are uppercase in the output string - All other characters are unchanged ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ //Two more semantically similar few-shots example TASKS ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ TASK: Write a method in Dafny to find maximum run of uppercase characters in the given string.
@@
SYSTEM: You are an expert AI assistant that writes Dafny programs. You are very good at writing verifiable correct code in terms of preconditions and postconditions of methods, and at finding the appropriate loop invariants for the pre/postconditions to hold. ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ TASK: Write a method in Dafny that matches a word containing ’z’. AI ASSISTANT: Step 1: Determine the method’s signature. In this case: method ContainsZ(s: string) returns (result: bool ) Step 2: Determine the preconditions. In this case: - There are no preconditions, the method will work. Strings and sequences are always not null. Step 3: Determine the postconditions. In this case: - The result is true if the input string contains the character ’z’ or ’Z’ - The result is false if the input string does not contain the character ’z’ or ’Z’ Step 4: Write the code, its formal specification, and the verification annotations that help the solver prove the postconditions. In this case: ‘‘‘dafny     method ContainsZ(s: string) returns (result: bool)         ensures result <==> (exists i :: 0 <= i < |s| && (s[i] == ’z’ || s[i] == ’Z’))     {         result := false;         for i := 0 to |s|             invariant 0 <= i <= |s|             invariant result <==> (exists k :: 0 <= k < i && (s[k] == ’z’ || s[k] == ’Z’))         {             if s[i] == ’z’ || s[i] == ’Z’ {                 result := true;                 break;             }         }     } ‘‘‘ ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ //Four more semantically similar few-shot example TASKS ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ TASK: Write a method in Dafny to find maximum run of uppercase characters in the given string. AI ASSISTANT:
Figure 4. An example of Specification Prompt (top) and CoT Prompt (bottom)
@@
In the next phase, we search for semantically similar Dafny methods in our MBPP-DFY-50 dataset, incorporating the LLM generated specification response. To perform the semantic similarity, we utilized OpenAI Embedding API that computes semantic similarity based on cosine distance. For the given specification generated in the first phase, we selected top five Dafny methods examples from our MBPP-DFY-50 dataset. We used 5 examples so that we can fit the whole prompt in LLMs’ context length (i.g., token limits). With these few-shot examples, we then prepared a CoT prompt that first instruct LLMs to synthesize the specifications according the problem decomposition steps and then synthesize a Dafny method that satisfies these specifications. Figure 4 (bottom) depicts an example of our designed CoT prompt.
@@ 3.5. Evaluation Metrics
We evaluate LLMs’ synthesized verified methods against the Dafny verifier. The verifier reports whether Dafny code is (a) syntactically correct and (b) verified. We report the synthesis performance using the pass@k metric (Kulal et al., 2019; Mark Chen et al., 2021; OpenAI, 2023a; Wang et al., 2023). In our study, we use verify@k metric (similar to pass@k), where k sample methods are synthesized per problem; a problem is considered to be verified if any attempt verifies.

### v2

@@ 3. Study Design
  • RQ2 [Signature Prompting]: How does additional context (method signature and tests) affect the synthesis outcome of LLMs?
  • RQ3 [Dynamic Few-Shot Prompting]: Do Chain of Thought (CoT) prompts supported by semantically similar few-shot examples and step-by-step problem decomposition improve the synthesis outcomes of LLMs?
@@ 3.1. Test Dataset: MBPP-san-DFY
To conduct our study, we needed a collection of problems with natural language descriptions and corresponding formally verified Dafny code. A search for Dafny code in GitHub yielded only 188 repositories, none of which had natural language specifications of the code. To use that code, we would have had to regenerate the informal natural language specifications from the formal code.
Instead, we adopted the Mostly Basic Programming Problems (MBPP) benchmark suite (Austin et al., 2021). MBPP is a collection of around 1,000 crowd-sourced algorithmic tasks with text description, function signature, concise implementation in Python, and three test cases. The MBPP tasks focus on basic arithmetic, array and integer sequences, and string processing. The textual descriptions of these tasks are designed to be concrete enough that an introductory Python programmer should be able to translate them into code, without any additional clarifications. Solutions to 427 of the MBPP tasks have been inspected manually, and published as a subset called MBPP-sanitized (MBPP-san). We based our data on MBPP-san. Our goal is to synthesize Dafny programs, rather than Python, so we removed 82 tasks that relied on Python-specific data types or external libraries [^4] Dafny does not have a standard library, much less external libraries. Note that any Dafny libraries would need to be formally-verified prior, or else the Dafny tasks would become much more complex than their Python versions. such as regex, math, heapq, etc., leaving a dataset of 345 problems.
In order to keep our study’s costs under budget, we furthered narrowed our focus to a randomly selected subset of 228 of these 345 problems employing sampling with 99% confidence interval and 5% margin of error. Finally we randomly picked a further subset of 50 tasks to act as verified Dafny exemplars, leaving us with 178 remaining problems for testing. We manually translated each of those 178 problem descriptions, signatures, and test cases into Dafny, but we did not translate the solutions’ Python code, or attempt to write specifications. This resulted in our main test dataset, MBPP-san-DFY: 178 instances of MBPP-san problems, each one with a description, method signature, and three test cases translated to suit Dafny.
The MBPP problem set and Python solutions to the MBPP problems are publicly available, and it seems likely they would have been incorporated into the LLMs’ training sets. There are no Dafny implementations of MBPP, however, so our MBPP-san-DFY dataset should not increase data leakage within our experiment.
@@ 3.2. Human Written Dataset: MBPP-DFY-50
To perform Dynamic Few-Shot prompting, we needed a relatively large and varied collection of verified Dafny methods that could serve as exemplars to be included in prompts. A search for an existing suitable collection proved fruitless. Our best alternative was to translate a subset of MBPP-san into Dafny ourselves, but now also translating the Python solution for each problem into Dafny, then coming up with specifications (method pre- and postconditions, loop invariants) for each example, and finally doing whatever annotation, correction, or rewriting was needed until the Dafny verifier would signal that the method was correct.
@@
3.3. Pilot Study and LLM Selection
Program synthesis in a verification-aware language requires knowledge of the target language syntax and semantics, plus the language implementation’s reasoning and theorem-proving abilities (Noble et al., 2022; First et al., 2023). To analyze LLMs’ reasoning capability in Dafny, we first performed a pilot study using our handwritten dataset MBPP-DFY-50. We started by looking at GPT-4 (OpenAI, 2023a) and GPT-3.5 (Mark Chen et al., 2021) from OpenAI (OpenAI, 2023b), PaLM-2 (Rohan Anil et al., 2023) and PaLM (Aakanksha Chowdhery et al., 2022) from Google, and CodeT5+ (Wang et al., 2021)and CodeT5 (Wang et al., 2023) from Salesforce, LLaMA and Code-LLaMA (Rozière et al., 2023) from Meta. All these LLMs have been explored in prior code generation, program synthesis, and reasoning-related studies.
Our pilot study consisted of querying all these LLMs with the problem description, asking each LLM to solve each problem. We then analyzed the LLM’s responses to get a sense of their reasoning abilities, and the quality of their approaches compared to our ground truth. For these 50 problems, CodeT5, CodeT5+, LLaMA, Code-LLaMA, and PaLM were unable to synthesize a single syntactically correct Dafny method with specifications. Even GPT-3.5 could generate syntactically correct solutions for 5 of the 50 pilot study problems — although GPT-4 and PaLM-2 could generate verified Dafny methods in 16/50 and 7/50 problems respectively. Following these results, we decided to focus our investigation onto GPT-4 and PaLM-2 only.
Contextless Prompt You are an expert Dafny programmer, and here is your task:
Write a method in Dafny to find maximum run of
uppercase characters in the given string.
You must return the method in the following form:
```dafny //Dafny Code ```
Signature Prompt You are an expert Dafny programmer, and here is your task:
Write a method in Dafny to find maximum run of uppercase
characters in the given string.
Method Signature: method maxRunUppercase (str: string ) returns (res: int ) Your code should satisfy these tests: var out1:=maxRunUppercase( "GeMKSForGERksISBESt" ); assert out1==5; var out2:=maxRunUppercase( "PrECIOusMOVemENTSYT" ); assert out2==6; var out3:=maxRunUppercase( "GooGLEFluTTER" ); assert out3==4; You must return the method in the following form: ```dafny //Dafny Code ```
Figure 2. Examples of Contextless and Signature prompts
@@ 3.4. Prompt Design
Based on the background literature and our pilot study, we knew that our experiment would depend on the prompts we offered to the LLMs. We experimented with three different prompts:
3.4.1. RQ1 [Contextless Prompt]:
@@
To answer RQ1 we designed a simple Contextless prompt that identifies the programming language (Dafny), the problem description, and formatting instructions for the code response. Our prompt design is inspired by Austin et al. (2021), and works well for popular languages like Python, JavaScript, Java, C#, and C++. The Contextless prompt instructs LLMs to synthesize a Dafny method in a single step without any additional context. Figure 2 (top) shows an example of a Contextless prompt. Writing implementation code in Dafny is similar to writing code in other high-level programming languages, so our goal with Contextless prompt is to establish a baseline of the LLMs’ performance in Dafny method synthesis.
3.4.2. RQ2 [Signature Prompt]:
@@
For RQ2, we designed Signature prompt with the required method signature and three test cases (see Figure 2, bottom) that extends Contextless prompt. Code generation experiments for non-verified code often perform better when prompted with signature and test cases (Austin et al., 2021; Mark Chen et al., 2021); we hypothesized that adding method signature and test cases would also benefit LLMs generating code that will be formally verified.
3.4.3. RQ3 [Dynamic Few-Shot Prompt]:
Chain of Thought (CoT) reasoning, which decomposes a problem into multiple intermediate steps, can lead to drastically improved performance for code synthesis tasks (Madaan et al., 2022; Wei et al., 2022; Reynolds and McDonell, 2021; Beurer-Kellner et al., 2023). Combined with relevant examples (aka few-shot prompting), we can prompt LLMs to generate solutions to problems that have little coverage in the training data. Our Chain of Though (CoT) prompts comprise five examples of tasks and solutions, where each example decomposes the problem step-by-step, and explains how to derive the implementation in Dafny. The rest of this subsection describes in detail how we engineer CoT prompts: Figure 4 (bottom) depicts one resulting CoT prompt while Figure 4 (top) shows an intermediate specification prompt, produced as one step in this process.
The design of our CoT prompt was informed by our own experience with Dafny (Noble et al., 2022; Noble, 2024). When we write a Dafny method, we start by defining the method’s signature, followed by thinking about the pre- and postconditions that should hold. We then go on to write the method body code, implementing algorithms that should produce the expected results. Finally we address validity, experimenting with assertions, loop invariants, etc., until the Dafny verifier is able to prove that the program is correct.
Finding examples with Retrieval-Augmented Generation (RAG)
For a given problem, a prompt will be most effective when it includes semantically similar exemplars. While creating our MBPP-DFY-50 dataset, we observed that the postconditions are the most critical parts of Dafny method specifications, and we identified a pattern: methods having similar signatures and postconditions require similar Dafny solutions.
Given that we were already using LLMs to find solutions, we were also able to employ the same LLMs to find prompts, a technique known as prompt chaining (Wei et al., 2022) or model cascading (Dohan et al., 2022). For each new problem, in a first phase we prompt the LLM to generate method specifications (signature and postconditions). Then in a second phase, we used the results of the first phase to prompt for Dafny method implementations. The inclusion of semantically-similar context in prompts is known as Retrieval Augmented Generation (RAG) (Lewis et al., 2020). Figure 3 demonstrates the resulting high-level architecture of our RAG pipeline for prompt chaining.
@@ 3.4.3. RQ3 [Dynamic Few-Shot Prompt]:
Figure 3. A high-level architecture of the RAG prompt-chaining pipeline for Dynamic Few-Shot Prompting.
@@
For a given problem, in the first phase we prompt the LLM to generate the method specifications —signature and postconditions. To teach LLM how to generate such specifications, we provide three examples of problem descriptions, method signatures, and postconditions, taken from our MBPP-DFY-50 dataset. We select these examples based on the semantic similarity of the problem descriptions. Figure 4 (top) shows an example of a first phase specification prompt.
@@
Dynamic Few-Shot Prompting [Specification Prompt] TASK: Write a method in Dafny to convert a given string to uppercase. Method Signature: method ToUppercase(s: string ) returns (v: string ) Postconditions: - The length of the returning string must be the same as the length of the input string - All lowercase characters of the input string are uppercase in the output string - All other characters are unchanged ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ //Two more semantically similar example TASKS ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ TASK: Write a method in Dafny to find maximum run of uppercase characters in the given string.
@@
Dynamic Few-Shot Prompting [Chain of Thought (CoT) Prompt] SYSTEM: You are an expert AI assistant that writes Dafny programs. You are very good at writing verifiable correct code in terms of preconditions and postconditions of methods, and at finding the appropriate loop invariants for the pre/postconditions to hold. ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ TASK: Write a method in Dafny that matches a word containing ’z’. AI ASSISTANT: Step 1: Determine the method’s signature. In this case: method ContainsZ(s: string) returns (result: bool ) Step 2: Determine the preconditions. In this case: - There are no preconditions, the method will work. Strings and sequences are always not null. Step 3: Determine the postconditions. In this case: - The result is true if the input string contains the character ’z’ or ’Z’ - The result is false if the input string does not contain the character ’z’ or ’Z’ Step 4: Write the code, its formal specification, and the verification annotations that help the solver prove the postconditions. In this case: ‘‘‘dafny method ContainsZ(s: string) returns (result: bool) ensures result <==> (exists i :: 0 <= i < |s| && (s[i] == ’z’ || s[i] == ’Z’)) { result := false; for i := 0 to |s| invariant 0 <= i <= |s| invariant result <==> (exists k :: 0 <= k < i && (s[k] == ’z’ || s[k] == ’Z’)) { if s[i] == ’z’ || s[i] == ’Z’ { result := true; break; } } } ‘‘‘ ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ //Four more semantically similar example TASKS ------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------ TASK: Write a method in Dafny to find maximum run of uppercase characters in the given string. AI ASSISTANT:
Figure 4. Examples of Specification Prompt (top) and CoT Prompt (bottom)
@@
In the second phase, we search for semantically similar Dafny methods in our MBPP-DFY-50 dataset, incorporating the LLM’s first phase response. We use the OpenAI embedding API to compute semantic similarity based on cosine distance. For the specification generated in the first phase, we select the top five Dafny similar examples from our MBPP-DFY-50 dataset. We use five examples so that the whole prompt fits inside the LLMs’ context length. With these examples, we then prepare a CoT prompt that first instructs the LLM to synthesize the specifications according the problem decomposition steps, and then synthesize a Dafny method that satisfies these specifications — see Figure 4 (bottom).
@@ 3.5. Evaluation Metrics
We evaluate LLMs’ synthesized methods using the Dafny verifier. The verifier reports whether Dafny code is (a) syntactically correct and (b) verified. We report the synthesis performance using the pass@k metric (Kulal et al., 2019; Mark Chen et al., 2021; OpenAI, 2023a; Wang et al., 2023). In our study, we use verify@k metric (similar to pass@k), where k sample methods are synthesized per problem; a problem is considered to be verified if any attempt verifies.